OpenStudy (anonymous):

Calculate the discriminant and use it to determine how many real-number roots the equation has. 3x2 – 6x + 1 = 0 A. one real-number root B. three real-number roots C. two real-number roots D. no real-number roots

OpenStudy (anonymous):

Compare the given equation with ax^2 + bx + c = 0: a = 3, b = -6, c = 1 Discriminant, D = b^2 - 4ac = (-6)^2 - 4 (3)(1) = 36 - 12 = 24 > 0 Two real roots.